The Taylor Company has an ROA of 7.6 percent, a profit margin of 5.2 percent, and an ROE of 14 percent.


a. What is the company's total asset turnover? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

b. What is the equity multiplier? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

a) Calculation of asset turnover ratio:

Asset turnover ratio= return on assets/profit margin

Asset turnover ratio= 7.6/5.2= 1.46

Asset turnover ratio= 1.46 times

b) Calculation of equity multiplier:

Equity multiplier= ROE/ROA

Equity multiplier= 14/7.6= 1.84

Equity multiplier= 1.84